A block is sliding without friction down a ramp along with a ring, a disc and a solid sphere. The height of the ramp is 26 meters.
What is the speed of the ring?
m/s

Tries 0/2

What is the speed of the disc?
m/s

Tries 0/2

What is the speed of the solid sphere?
m/s

Tries 0/2

What is the speed of the block at the bottom of the ramp?
Answer #1

solution -10 R2 -16 =mv2 (a) Mgn = 12 mv2 + 1/ 1 W2 (Forring) h = MV2+ mo?xV2 Rolling motion = mv2t2 mv2 v²= gh. or v=Jgh OR